OpenStudy (anonymous):

Ten persons compete in a race. In how many ways can the 1st, 2nd, and 3rd place awards be distributed, if only one person receives each award?

OpenStudy (anonymous):

1st: Possible Outcomes: 10 2nd: Possible Outcomes: 9 (as 1st is taken by 1 person, so 10-1 =9) 3rd: Possible Outcomes: 8 (1st and 2nd are taken, so 10-2=8) Not just multiply 10*9*8 = 720

OpenStudy (anonymous):

I agree with KriminalDrama